OpenStudy (anonymous):

3 1/3 x 3/8= Not Here 1 2/8 1 4/24 1 1/4 3 3/24

sam (.sam.):

1 1/4

sam (.sam.):

\[3\frac{1}{3}X\frac{3}{8}\] \[\frac{10}{3}X\frac{3}{8}\] Common denominator \[\frac{5}{4}=1\frac{1}{4}\]

Directrix (directrix):

3 1/3 x 3/8= 10/3 * 3/8 = 10/8 = 5/4 = 1 and 1/4

Directrix (directrix):

@ airChick --> Were you to give the answer in simplest form? If not, that first option of 1 and 2/8 simplifies to 1 and 1/4 which is also the third option.